In a factory, a parabolic mirror to be used in a searchlight was placed on the floor. It measured 50 centimeters tall and 90 centimeters wide. Find the equation of the parabola.
Answer by lwsshak3(11628)

In a factory, a parabolic mirror to be used in a searchlight was placed on the floor. It measured 50 centimeters tall and 90 centimeters wide. Find the equation of the parabola.
***
equation of a parabola that opens up with vertex at the origin: x^2=4py
point on the parabola:(45,50)(top right edge of parabolic lens)
equation: 45^2=4p(50)
4p=45^2/50=2025/50=40.5
equation:
x^2=(40.5)y